OpenStudy (anonymous):

For the following system, use the second equation to make a substitution for y in the first equation. 3x + y = 2 y + 4 = 3x 3x + 3x - 4 = 2 3x + 3 - 4x = 2 3x + 3x + 4 = 2

OpenStudy (anonymous):

3x+3x-4=2

OpenStudy (anonymous):

x + y = 4 x - 2y = -5 If the first equation is multiplied by 2 and then the equations are added, the result is _____. 3x = 13 3x = 3 x = 3

OpenStudy (anonymous):

3x=3

OpenStudy (anonymous):

awesome thnx

OpenStudy (anonymous):

Which of the systems below is equivalent to the given system? 0.01x - 0.05y = 0.1 0.4x + 0.3y = 0.9 10x - 5y = 10 and 4x + 3y = 9 x - 5y = 10 and 4x + 3y = 9 x - 5y = 1 and 4x + 3y = 9

OpenStudy (anonymous):

the second one.

OpenStudy (anonymous):

Given the system of equations, what is the solution? 2x + y = -1 x - y = -5 {(-8, -3)} {(2, 7)} {(-2, 3)}

OpenStudy (anonymous):

#3

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

Simplify. 6a - {b - [3a - (2b + c) + 4a - (a + 2b - c)]}. 12a - 5b 6a - b + 2c 12a - 5b + 2c b

OpenStudy (anonymous):

number 1) 12a-5b